A reaction has the rate law rate = k[A][B].

Calculate the rate constant k if:

[A] = 0.15 mol L-1

[B] = 0.20 mol L-1

rate = 3.11 x 10-2 mol L-1 s-1
1
Expert's answer
2019-03-22T04:28:41-0400

r = k[A][B]

k = r / ([A][B]) = 3.11 × 10-2 mol×L-1×s-1 / (0.15 mol×L-1 × 0.20 mol×L-1) = 1.0367 L×mol-1×s-1 ≈ 1.04 L×mol-1×s-1.